# Decoherence of electron spin qubit during transfer between two semiconductor quantum dots at low magnetic fields

### v2

@@ III.3 Spin dephasing $\langle\delta\varphi^{2}\rangle$
As a last source of error during coherent transfer we discuss pure dephasing effects, caused by the uncontrolled, random evolution of the phase of electron qubit during the transition. Main source of such randomness is non-zero dot-dependent Zeeman splitting combined with a random time spend in the dots. It is caused by a presence of both low- and high- frequency fluctuations of electric field as well as the phonon induced transitions, and can be parameterized by a random excess of time spent in one of the dots $\delta t$. It translates into stochastic contribution to a relative phase between the spin eigenstates:
@@
For a symmetric sweep the electron on average spends the same amount of time in each dot, i.e. $\langle\delta t\rangle=0$ and dephasing is related to non-zero $\langle\delta t^{2}\rangle$. Note that even if a given mechanism leads to finite $\langle\delta t\rangle$, this simply leads to renormalization of the deterministic phase acquired by the spin during the shuttling, while the presence of nonzero $\langle\delta t^{2}\rangle-\langle\delta t\rangle^{2}$ leads to dephasing. The gradient of Zeeman splittings $\Delta E_{Z}$ often originates from non-zero difference in g-factors between the two dots, i.e. $\Delta E_{Z}=(\delta g/g)E_{Z}$, or alternatively the intentional gradient of magnetic field. In both cases we expect $\Delta E_{Z}\ll 2t_{c}$, which allows to treat $\Delta E_{Z}$ as small perturbation.
III.3.1 Non-adiabatic dephasing
As a first relevant mechanism leading to $\langle\delta t^{2}\rangle\neq 0$, we consider the dephasing during charge relaxation between the eigenstates of Fig. 1 b). They are caused by the energy exchange between the electron and the environment, that is in resonance with the instantaneous orbital gap $\Omega_{s}(t)$ from Eq. (3). In the limit of a single excitation at avoided crossing, the contribution $\langle\delta\varphi^{2}\rangle$ corresponds to a random period of time spent in the excited orbital state. This time period is set by the time between the excitation and subsequent relaxation, which brings the electron charge back to the ground state. To illustrate this mechanism, we set the excitation probability to $P_{e}$ and the relaxation rate to a constant $\Gamma_{-}$. In such a case the waiting time is given by the distribution $p(\delta t)=\Gamma_{-}e^{-\Gamma_{-}\delta t}$, with a variance $\langle\delta t^{2}\rangle=\Gamma_{-}^{-2}$. Directly from the above we have [27]:
@@ III.3.1 Non-adiabatic dephasing
In Fig. (4), we use dots to compare the results of AME (Eq.(11)) with approximate expression (Eq. (25) for GaAs (a) and Si (b) at different $\Delta E_{Z}$ values. The simple model assumes constant relaxation rate and spin-splitting, likely overestimating AME due to relaxation and effectively lower $\Delta E_{Z}$ around $\epsilon\ll 2t_{c}$. In GaAs, negligible dephasing at $\Delta E_{Z}\!\leqslant\!0.1$ $\mu$eV is attributed to fast relaxation with strong phonon coupling. Si exhibits lower relaxation rates, leading to longer times in the excited state. For $\Delta E_{Z}\!=\!0.1\,\mu$eV (orange), the error combined with losing charge ($Q_{\text{charge}})$ has a single minimum, indicating significant phase loss during charge relaxation. Fig. 4 reveals a region where dephasing during relaxation is effective inactive, and the error is dominated by $Q_{\text{charge}}$. This is no longer true at slower sweeps, where the dashed lines align quantitatively with AME at $v\!<\!50$ $\mu$eV/ns for GaAs and $v\!<\!10$ $\mu$eV/ns for Si.
Figure 4: Dephasing during inelastic tunneling due to random time spent in the excited state as a function of sweep rate $v$ for GaAs (a) and SiGe (b) devices, with relaxation rates from Appendix C. We show results for the tunnel coupling of $2t_{c}=40\mu$eV and for different values of Zeeman splittings difference $\Delta E_{Z}=10^{-1}\,\mu$eV, $1\,\mu$eV for GaAs and $\Delta E_{Z}=10^{-3}\,\mu$eV, $10^{-2}\,\mu$eV amd $10^{-1}\,\mu$eV in Si. Dots corresponds to solutions of Adiabatic Master Equation without averaging over slow noise, while the dashed lines lines represents approximate formula Eq. (24). For the case of Si and $\Delta E_{Z}=10^{-1}$ (green solid lines), the analytical formula overestimates the error since $\Delta E_{Z}/\Gamma_{-}\geq 1$.
III.3.2 Adiabatic dephasing
@@
$\Omega_{s}(t)=\Big{\{}\big{[}\epsilon+\delta\epsilon+\tfrac{\sigma_{s}}{2}(\Delta E_{Z}+$ $\delta E_{Z,L}-\delta E_{Z,R})\big{]}^{2}$
@@
$\varphi=\int_{0}^{t_{f}}\omega_{\text{spin}}(t)\,\text{d}s=\varphi_{0}+\delta\varphi,$ (28)
@@
which after averaging over realisations of all four quasistatic processes leads to dephasing $Q\approx\tfrac{1}{2}\langle\delta\varphi^{2}\rangle$.
@@
Figure 5: Total coherent transfer error $Q$ as a function of sweep rate $v$ for two values of tunnel couplings $2t_{c}=40\,\mu$eV (filled dots) and $2t_{c}=100\,\mu$eV (hollow dots), with results for a GaAs DQD in panel (a), and for a Si DQD in panel (b). Dots connected by lines represent solutions of the Adiabatic Master Equation averaged over quasistatic noises in spin splitting and detuning, while the dashed and dotted lines are the analytical predictions due to separate mechanisms. The pink lines are the depahsing error from Eq. (29) due to $1/f$ noise in detuning having an amplitude of $S_{1}=0.5$ $\mu$eV/$\sqrt{\mathrm{Hz}}$, marked by the corresponding $\Delta E_{Z}$, and the blue lines are phase error due to spin-splitting fluctuations in two QDs marked by corresponding $T_{2}^{*}$ values, Eq. (30). For Si we also plot the phase error due to charge transfer error caused by high-frequency noise, Eq. (24), as gray dashed lines, corresponding to two values of $2t_{c}$. In both panels the colors of solid lines correspond to distinct values of spin-orbit couplings (SOCs). For GaAs in (a) we plot the results for intrinsic SOC $|a|=1\,\mu$eV in red, and for synthetic SOC $b=1\,\mu$eV ($a\ll 1\mu$eV) in black. In the numerical simulation of the full error we use use $T_{2}^{*}\!=\!2$ $\mu$s and $\Delta E_{Z}=10^{-1}\mu$eV. For Si (b) we have the intrinsic SOC $|a|=0.1\mu$eV, and we compare the AME results for synthetic SOC $b\!=\!0$ (green) versus $b=1\,\mu$eV (yellow), $T_{2}^{*}=20\,\mu$eV and $\Delta E_{Z}=10^{-3}\,\mu$eV. Influence of dephasing due to $1/f$ noise in detuning for larger values of $\Delta E_{Z}$ on the total error can be estimated by taking a maximum of solid line and the pink dashed line corresponding to the relevant $\Delta E_{Z}$. The remaining parameters can be found in Table 1.
@@
Calculation of $\delta\varphi$ to the lowest order in fluctuating quantitites $\delta\epsilon$, $\delta E_{Z,L(R)}$, and $\delta t_{c}$ is given in Appendix D. Let us now discuss the results for $\langle\delta\varphi^{2}\rangle$ in two limits, in which it is has a simple and physically transparent form. For sufficiently large $\Delta E_{Z}$ the most relevant mechanism leading to non-zero $\langle\delta\varphi^{2}\rangle$ are the fluctuations of dots energy detuning $\delta\epsilon$, which are directly related to the electrical noise and can be seen as a uncertainty in the exact position of the avoided crossing. In such a case similarity to the previous section, a time period spent in each dot is random, and hence $\langle\delta t^{2}\rangle$ becomes non-zero (See Eq. 23). The value of $\delta\epsilon$ shifts initial and final values of the detunings ±$\pm vt_{f}+\delta\epsilon$, which directly translates into $\delta t\approx\delta\epsilon/v$, and hence:
@@
Note that due to symmetry reasons slow, quasistatic fluctuations in tunnel coupling $\delta t_{c}$ are not expected to introduce any significant random spin phase $\delta\varphi$. For more quantitative derivation of Eq. (29) and effects of tunnel coupling noise see Appendix D. We highlight that for typical amplitude of those fluctuations $\delta t_{c}\ll\delta\epsilon$, and in the limit of $v\ll t_{c}^{2}$ significant modification of the coherent charge excitations $Q_{\text{LZ}}$ are not seen in the numerical results (See [53] for this effect).
In the limit of sufficiently low $\Delta E_{Z}$ the dephasing is dominated by a standard inhomogenous broadening, caused by the fluctuations of spin-splitting. Assuming fluctuations in the dots are independent $\langle\delta E_{Z,L}\delta E_{Z,R}\rangle\!=\!0$ and $\langle E_{Z,L}^{2}\rangle=\langle\delta E_{Z,R}^{2}\rangle=\sqrt{2}/T_{2}^{*}$, the contribution to a random phase reads:
@@
For numerical test of approximate analytical formulas Eqs. (29) and (30) see Appendix  D.
@@ IV Case study: GaAs and Si DQD devices
We conclude this paper with the case study of the coherent double-dot transition in isotopically purified Si ($T_{2}^{*}=20\mu$s) and compare it against GaAs with continuous estimation of the Overhauser field [41, 42] ($T_{2}^{*}=2\mu$s). The remaining parameters of the devices are discussed in Section II.4. The main result of the paper is shown in Fig. 5, where we the dots connected with solid lines mark the coherent transfer error of equal spin superposition computed using the Adiabatic Master Equation and averaged over low-frequency fluctuations (see Sec. II). For the two DQD devices, we use transition rates from Appendix C and compute the transfer error $Q$ for the combination of parameters $2t_{c}=40,100\,\mu$eV (filled and hollow dots respectively), and a selection of relevant parameters: single-dot spin coherence time $T_{2}^{*}=2\,\mu$s (GaAs with field narrowing), $T_{2}^{*}=20\,\mu$s (Si), spin splitting gradient $\Delta E_{Z}$ between $10^{-3}$ and $1$ $\mu$eV and various spin-orbit couplings. For GaAs, we compare the relatively strong intrinsic $|a|=1\,\mu$eV (red) against spin-orbit coupling originating from the magnetic field gradient close to the manipulation region $b_{\perp}=1\mu$eV (black). For Si, we use the same value of synthetic SOC (yellow) but use a smaller intrinsic $|a|=0.1\,\mu$eV (green). We additionally add lines corresponding to analytical predictions, dephasing noise related to spin splitting (blue, dashed), dephasing due to detuning fluctuations (violet dashed), and charge excitation using single excitation approximation limit (gray dashed). For comparison, we also add the line corresponding to $T_{2}^{*}=20$ns without field narrowing in GaAs (dotted blue line in a).
@@
As expected, in the limit of large sweep rates, in both devices transfer error $Q$ is limited by charge excitation due to the Landau-Zener mechanism. Starting from the largest considered sweep rates, decreasing $v$ is initially improving the transfer until another mechanism becomes dominant. Those mechanisms, contrary to LZ, typically worsen the transfer as the transfer time increases, i.e. as $v$ decrease. This gives rise to an interesting trade-off mechanism we discuss below.
@@
In GaAs, the subsequent mechanism following the reduction of $v$ is spin relaxation induced by spin-orbit mixing near avoided crossings. Despite using the same value of intrinsic and synthetic spin-orbit coupling (when the latter is finite), a difference between the influence of the two couplings is visible, in particular at large $2t_{c}=100\,\mu$eV, for which the intrinsic one suffers from Van Vleck cancellation since $2t_{c}\gg E_{Z}=30\mu$eV (Red, hollow dots in the figure). The best achievable in this scenario $Q=10^{-5}$ takes place at $v=10^{3}\,\mu$eV/ns and increases by the order of magnitude if the SOC is of the intrinsic character. For smaller $2t_{c}=40\,\mu$eV one can achieve $Q=5\times 10^{-3}$ around $v=10^{3}\,\mu$eV, with no visible difference between the two SOCs. At slower sweeps the $Q$ is further increasing, until the error becomes dominated by the charge noise (dashed violet line), which changes the scaling to $Q\propto 1/v^{2}$.
@@
Alternatively, at the slowest sweeps the error can be dominated by the spin splitting fluctuations in the dots. The position at which it happens depends on $T_{2}^{*}$ (see the dashed blue line for $T_{2}^{*}=2\,\mu$s). It means that for GaAs without field monitoring ($T_{2}^{*}=20$ns) this mechanism can become the dominant (see dotted blue line), and effectively create a single minimum of $Q=10^{-4}$ at $2t_{c}=100\,\mu$eV and $Q=10^{-2}$ at $2t_{c}=40\,\mu$eV.
@@
In Si, a similar behaviour of a trade-off, between Landau-Zener excitations and a mechanism of dephasing due to low-frequency noise, is expected at $\Delta E_{Z}=1\mu$eV (pink dotted line in b). It means that close to large longitudinal gradients (manipulation regions) the error might be not smaller than $Q=10^{-5}$ at $2t_{c}=100\,\mu$eV and $Q=10^{-4}$ at $2t_{c}=40\,\mu$eV. Away from that regions, where order of magnitude-wise $\Delta E_{Z}=10^{-3}$, the shape of $Q$ is more complex.
@@
The spin relaxation mechanism in Si becomes relevant only in regions with a large transverse gradient (depicted by yellow lines with $b_{\perp}=1,\mu$eV). In its presence, especially at larger $2t_{c}=100,\mu$eV, the shape of the $Q$ curve (the yellow line with hollow dots) resembles the GaAs case, exhibiting a transition from Landau-Zener (L-Z) dominance at large $v$ to spin relaxation dominance at small $v$. For all the other parameter sets, a transition occurs to regime dominated by incoherent charge transfer error that is well described in the SEAL approximation, marked by gray dashed lines. The interplay between these two charge transfer errors results in a minimum, reaching as low as $Q=10^{-7}$ at $v=10^{3},\mu$eV/ns with a larger tunnel coupling of $2t_{c}=100\,\mu$eV, and $Q=10^{-4}$ at $v=200,\mu$eV/ns with smaller $2t_{c}=40\,\mu$eV.
@@
V Discussion and summary
We now discuss implications of our results for coherent spin transfer through a long ($\sim 1-10$ $\mu$m) quantum dot array. For the mechanism of charge transfer error and spin relaxation, $Q$ is additive: the error accumulated during transfer is a sum of the contributions from each dot-to-dot transfers, i.e.
$Q_{N}=\sum_{n=1}^{N}Q_{\text{charge}}^{(n)}+Q_{\text{relax}}^{(n)},$ (31)

### v3

@@ III.3 Spin dephasing $\langle\delta\varphi^{2}\rangle$
As the last source of error during coherent transfer we discuss pure dephasing effects, caused by the uncontrolled, random evolution of the phase of electron qubit during the transition. This randomness appears when $\Delta E_{Z}$ is nonzero due to dot-dependent Zeeman splitting, and when the time spent in each of the dots is a random variable. This randomness is caused by incoherent transitions between instantaneous eigenstates due to presence of both low- and high- frequency fluctuations of electric field, as well as electron-phonon coupling. It can be parameterized by a random excess of time spent in one of the dots, $\delta t$, which translates into stochastic contribution to a relative phase between the spin eigenstates:
@@
In the fully adiabatic case, and for a symmetric sweep, the electron on average spends the same amount of time in each dot, i.e. $\langle\delta t\rangle=0$ and dephasing is related to non-zero $\langle\delta t^{2}\rangle$. Note that even if a given mechanism leads to finite $\langle\delta t\rangle$, this simply gives renormalization of the deterministic phase acquired by the spin during the shuttling, while the presence of nonzero $\langle\delta t^{2}\rangle-\langle\delta t\rangle^{2}$ leads to dephasing. In general, we expect $\Delta E_{Z}\ll 2t_{c}$, which allows to treat $\Delta E_{Z}$ as small perturbation.
III.3.1 Dephasing due to high-frequency noise
As a first relevant mechanism leading to $\langle\delta t^{2}\rangle\neq 0$, we consider the dephasing caused by high-frequency-noise induced transitions between the eigenstates of Fig. 1 b). They are caused by the energy exchange between the electron and the environment, that is in resonance with the instantaneous orbital gap $\Omega_{s}(t)$ from Eq. (3). In the limit of a single excitation at avoided crossing, the contribution $\langle\delta\varphi^{2}\rangle$ is due to variance of distribution of time spent in the excited orbital state. This time period is set by the time between the excitation and subsequent relaxation, which brings the electron charge back to the ground state. To illustrate this mechanism, we set the excitation probability to $P_{e}$ and the relaxation rate to a constant $\Gamma_{-}$. In such a case the waiting time is given by the distribution $p(\delta t)=\Gamma_{-}e^{-\Gamma_{-}\delta t}$, with a variance $\langle\delta t^{2}\rangle=\Gamma_{-}^{-2}$. Directly from the above we have [33]:
@@ III.3.1 Dephasing due to high-frequency noise
In Fig. (4), we use dots to compare the results of AME (Eq.(11)) with approximate expression (Eq. (25) for GaAs (a) and Si (b) at different $\Delta E_{Z}$ values. The simple model assumes constant relaxation rate and spin-splitting. It is likely to overestimate the error calculated using the AME, where $\Delta E_{Z}$ is effectively lower around avoided crossing, i.e. when $\epsilon\ll 2t_{c}$. In GaAs, negligible dephasing at $\Delta E_{Z}\!\leqslant\!0.1$ $\mu$eV is attributed to fast relaxation with strong electron-phonon coupling. Si exhibits lower relaxation rates, leading to longer times in the excited state. For $\Delta E_{Z}\!=\!0.1\,\mu$eV (orange), the pure dephasing error combined with charge transfer error ($Q_{\text{charge}})$ has a single minimum, indicating significant phase loss during charge relaxation. Fig. 4 reveals a region where dephasing during relaxation is ineffective, and the error is dominated by $Q_{\text{charge}}$. This is no longer true at slower sweeps, where the dashed lines align quantitatively with AME at $v\!<\!50$ $\mu$eV/ns for GaAs and $v\!<\!10$ $\mu$eV/ns for Si.
Figure 4: The sum of charge transfer error $Q_{\text{charge}}$ and dephasing during inelastic tunneling due to random time spent in the excited state $\tfrac{1}{2}\langle\delta\varphi^{2}\rangle_{\text{non-adiab}}$, as a function of sweep rate $v$ for GaAs (a) and SiGe (b) devices. We show results for the tunnel coupling of $2t_{c}=40\mu$eV and for different values of Zeeman splittings difference $\Delta E_{Z}=10^{-1}\,\mu$eV, $1\,\mu$eV for GaAs and $\Delta E_{Z}=10^{-3}\,\mu$eV, $10^{-2}\,\mu$eV and $10^{-1}\,\mu$eV in Si (colors). Dots corresponds to solutions of Adiabatic Master Equation without averaging over slow noise, while dashed lines represent the approximate formula from Eq. (24). While at $v\sim 10^{2}$ $\mu$eV/ns the error is dominated by charge transfer error analyzed in Sec. III.1, at lower $v$ dephasing during inelastic tunneling dominates, as the total error becomes dependent on the value of $\Delta E_{Z}$. For the case of Si and $\Delta E_{Z}=10^{-1}$ (yellow lines), the analytical formula overestimates the error since $\Delta E_{Z}/\Gamma_{-}\geq 1$. We use relaxation rates from Appendix C.
III.3.2 Dephasing due to low-frequency (quasistatic) noise
@@
$\Omega_{s}(t)=\Big{\{}\big{[}\epsilon(t)+\delta\epsilon+\tfrac{\sigma_{s}}{2}(\Delta E_{Z}+$ $\delta E_{Z,L}-\delta E_{Z,R})\big{]}^{2}$
@@
$\varphi=\int_{0}^{t_{f}}\omega_{\text{spin}}(t)\,\text{d}t=\varphi_{0}+\delta\varphi,$ (28)
@@
which after averaging over realisations of all four quasistatic processes leads to dephasing $Q_{\varphi}\approx\tfrac{1}{2}\langle\delta\varphi^{2}\rangle$.
@@
Figure 5: Total coherent transfer error $Q$ as a function of sweep rate $v$ for two values of tunnel couplings $2t_{c}=40\,\mu$eV (filled dots) and $2t_{c}=100\,\mu$eV (hollow dots), with results for a GaAs DQD in panel (a), and for a Si DQD in panel (b). Dots connected by lines represent solutions of the Adiabatic Master Equation averaged over quasistatic noises in spin splitting and detuning, while the dashed and dotted lines are the analytical predictions due to separate mechanisms. The violet lines are the depahsing error from Eq. (29) due to $1/f$ noise in detuning having an amplitude of $S_{1}=0.5$ $\mu$eV/$\sqrt{\mathrm{Hz}}$, marked by the corresponding $\Delta E_{Z}$, and the blue lines are phase error due to spin-splitting fluctuations in two QDs marked by corresponding $T_{2}^{*}$ values, Eq. (30). For Si we also plot the phase error due to charge transfer error caused by high-frequency noise, Eq. (24), as gray dashed lines, corresponding to two values of $2t_{c}$. In both panels the colors of solid lines correspond to distinct values of spin-orbit couplings (SOCs). For GaAs in (a) we plot the results for intrinsic SOC $|a|=1\,\mu$eV in red, and for synthetic SOC $b=1\,\mu$eV ($a\ll 1\mu$eV) in black. In the numerical simulation of the full error we use use $T_{2}^{*}\!=\!2$ $\mu$s and $\Delta E_{Z}=10^{-1}\mu$eV. For Si (b) we have the intrinsic SOC $|a|=0.1\mu$eV, and we compare the AME results for synthetic SOC $b_{\perp}\!=\!0$ (green) versus $b_{\perp}=1\,\mu$eV (yellow), $T_{2}^{*}=20\,\mu$eV and $\Delta E_{Z}=10^{-3}\,\mu$eV. Influence of dephasing due to $1/f$ noise in detuning for larger values of $\Delta E_{Z}$ on the total error can be estimated by taking a maximum of solid line and the violet dashed line corresponding to the relevant $\Delta E_{Z}$. The remaining parameters can be found in Table 1.
@@
Calculation of $\delta\varphi$ to the lowest order in fluctuating quantitites $\delta\epsilon$, $\delta E_{Z,L(R)}$, and $\delta t_{c}$ is given in Appendix D. Let us now discuss the results for $\langle\delta\varphi^{2}\rangle$ in two limits, in which it is has a simple and physically transparent form. For sufficiently large $\Delta E_{Z}$ the most relevant mechanism leading to non-zero $\langle\delta\varphi^{2}\rangle$ are the fluctuations of dots energy detuning $\delta\epsilon$, which are directly related to the electrical noise and can be seen as a uncertainty in the exact position of the avoided crossing. In such a case similarity to the previous section, a time period spent in each dot is random, and hence $\langle\delta t^{2}\rangle$ becomes non-zero (see Eq. 23). The value of $\delta\epsilon$ shifts initial and final values of the detuning ±$\pm vt_{f}+\delta\epsilon$, which directly translates into $\delta t\approx\delta\epsilon/v$, and hence:
@@
Note that due to symmetry reasons contribution from slow, quasistatic fluctuations in tunnel coupling $\delta t_{c}$ is vanishing in the leading order, and hence does not contribute significantly to $\delta\varphi$. For more quantitative derivation of Eq. (29) and effects of tunnel coupling noise see Appendix D. We highlight that for typical amplitude of those fluctuations $\delta t_{c}\ll\delta\epsilon$, and in the limit of $v\ll t_{c}^{2}$ significant modification of the coherent charge excitations $Q_{\text{LZ}}$ are not seen in the numerical results (see [67] for this effect).
In the limit of sufficiently low $\Delta E_{Z}$ the dephasing is dominated by standard inhomogenous broadening, caused by the fluctuations of spin-splitting. Assuming fluctuations in the dots are independent $\langle\delta E_{Z,L}\delta E_{Z,R}\rangle\!=\!0$ and $\langle E_{Z,L}^{2}\rangle=\langle\delta E_{Z,R}^{2}\rangle=\sqrt{2}/T_{2}^{*}$, the contribution to a random phase for a single transition reads:
@@
For numerical test of approximate analytical formulas, Eqs. (29) and (30), see Appendix  D.
@@ IV Case study: GaAs and Si DQD devices
We conclude this paper with the case study of the coherent double-dot transition in isotopically purified Si ($T_{2}^{*}\!=\!20$ $\mu$s) and compare it against GaAs with continuous estimation of the Overhauser field [53, 54] ($T_{2}^{*}\!=\!2$ $\mu$s). The remaining parameters of the devices are discussed in Section II.4. The main result of the paper is shown in Fig. 5, where the dots connected with solid lines mark the coherent transfer error of equal spin superposition, computed using the Adiabatic Master Equation and averaged over low-frequency fluctuations (see Sec. II). For the two DQD devices, we use transition rates from Appendix C and compute the transfer error $Q$ for the combination of parameters $2t_{c}=40,100\,\mu$eV (filled and hollow dots respectively), and a selection of relevant parameters: single-dot spin coherence time $T_{2}^{*}=2\,\mu$s (GaAs with field narrowing), $T_{2}^{*}=20\,\mu$s (Si), spin splitting gradient $\Delta E_{Z}$ between $10^{-3}$ and $1$ $\mu$eV and various spin-orbit couplings. For GaAs, we compare the relatively strong intrinsic $|a|=1\,\mu$eV (red) against spin-orbit coupling originating from the magnetic field gradient close to the manipulation region $b_{\perp}=1\mu$eV (black). For Si, we use the same value of synthetic SOC (yellow) but use a smaller intrinsic $|a|=0.1\,\mu$eV (green). We add lines corresponding to analytical predictions, dephasing noise related to spin splitting (blue, dashed), dephasing due to detuning fluctuations (violet dashed or dotted), and charge excitation using single excitation approximation limit (gray dashed). For comparison, we also add the line corresponding to $T_{2}^{*}=20$ ns without field narrowing in GaAs (dotted blue line in a).
@@
As expected, in the limit of large sweep rates $v$, in both devices transfer error $Q$ is limited by charge excitation due to the Landau-Zener mechanism. Starting from the largest considered sweep rates, decreasing $v$ is initially improving the transfer until another mechanism becomes dominant. Those mechanisms, contrary to LZ, typically worsen the transfer as the transfer time increases, i.e. as $v$ decrease. This gives rise to an interesting trade-off mechanism we discuss below.
@@
In GaAs, the mechanism that becomes dominant following the reduction of $v$ is spin relaxation induced by spin-orbit mixing near avoided crossings. Despite using the same value of intrinsic and synthetic spin-orbit coupling (when the latter is finite), a difference between the influence of the two couplings is visible, in particular at large $2t_{c}=100\,\mu$eV, for which the intrinsic one suffers from Van Vleck cancellation since $2t_{c}\gg E_{Z}=30\mu$eV (Red, hollow dots in the figure). The best achievable in this scenario $Q=10^{-5}$ takes place at $v=10^{3}\,\mu$eV/ns and increases by the order of magnitude if the SOC is of the intrinsic character. For smaller $2t_{c}=40\,\mu$eV one can achieve $Q=5\times 10^{-3}$ around $v=10^{3}\,\mu$eV, with no visible difference between the two SOCs. At slower sweeps the $Q$ is further increasing, until the error becomes dominated by the charge noise (dashed violet line), which changes the scaling to $Q\propto 1/v^{2}$.
@@
Alternatively, at the slowest sweeps the error can be dominated by the spin splitting fluctuations in the dots. The position at which it happens depends on $T_{2}^{*}$ (see the dashed blue line for $T_{2}^{*}=2\,\mu$s). It means that for GaAs without field monitoring ($T_{2}^{*}=20$ns) this mechanism becomes the dominant one (see dotted violet line), and effectively creates a single minimum of $Q(v)$, with values of $Q=10^{-4}$ at $2t_{c}=100\,\mu$eV , and $Q=10^{-2}$ at $2t_{c}=40\,\mu$eV.
@@
In Si, a similar trade-off between Landau-Zener excitations, and a mechanism of dephasing due to low-frequency noise, is expected at $\Delta E_{Z}=1\mu$eV (violet dotted line in b). It means that close to large longitudinal gradients (for DQD in the manipulation region) the error might not be smaller than $Q=10^{-5}$ at $2t_{c}=100\,\mu$eV and $Q=10^{-4}$ at $2t_{c}=40\,\mu$eV. Away from such region, where order of magnitude-wise $\Delta E_{Z}\!\sim\!10^{-3}$, the shape of $Q$ is more complex.
@@
The spin relaxation mechanism in Si becomes relevant only in regions with a large transverse gradient (depicted by yellow lines with $b_{\perp}=1,\mu$eV). In its presence, especially at larger $2t_{c}=100,\mu$eV, the shape of the $Q$ curve (the yellow line with hollow dots) resembles the GaAs case, exhibiting a transition from Landau-Zener (L-Z) dominance at large $v$ to spin relaxation dominance at small $v$. For all the other parameter sets, a transition occurs to regime dominated by incoherent charge transfer error that is well described in the SEAL approximation, marked by gray dashed lines. The interplay between these two charge transfer errors results in a minimum, reaching as low as $Q=10^{-7}$ at $v=10^{3}\,\mu$eV/ns with a larger tunnel coupling of $2t_{c}=100\,\mu$eV, and $Q=10^{-4}$ at $v=200\,\mu$eV/ns with smaller $2t_{c}=40\,\mu$eV.
@@
V Implications for transfer of a qubit through a chain of dots
We summarize the findings in Fig. 6, where we plot the predicted minimum of the coherent transfer error $Q$ as a function of tunnel coupling. For both devices, we maintain the device parameters as in Fig. 5, and we compare two cases: one where the transition occurs close to the manipulation region with $\Delta E_{Z}=b_{\perp}=1\,\mu eV$, and another where the transition is far from the manipulation region, in case of which we set the intentional gradient to zero. In this Section, the results shown in this Figure will be used to analyze the prospects for coherent spin transfer through a long ($\sim 1-10\,\mu$m) array of tunnel-coupled quantum dots.
Figure 6: Estimated lowest coherent transfer error $Q_{\text{min}}$ between neighouring dots, as a function of tunnel coupling, for the case of dots away from manipulation region in Si (green) and GaAs (red) quantum dots, compared against the same dots in the manipulation region depicted using yellow and black color, respectively. In the first case we use parameters from Table 1, and for manipulation region we set $b_{\perp}=\Delta E_{Z}=1\,\mu$eV. The dashed horizontal lines denotes the threshold for the single weak link, $Q\approx 10^{-3}$, and for the uniform array of $N=100$ dots, $Q=10^{-5}$.
We begin by discussing the mechanisms of charge transfer error and spin relaxation, which dominate the lowest achievable error $Q$, for the whole range of considered $2t_{c}>E_{Z}=30\mu$eV. For these processes, the total error accumulated during transfer is the sum of the contributions from each dot-to-dot transfer, i.e.,
$Q_{N}\approx\sum_{n=1}^{N}Q_{\text{charge}}^{(n)}+Q_{\text{relax}}^{(n)},$